OpenStudy (warpedkitten):

If the circulation for newspaper 1 is 384 thousand more than the circulation for newspaper​ 2, and their combined circulation is 2512 ​thousand, find the circulation for each newspaper.

OpenStudy (daniel_chernioglo):

@satellite73 this was a question posted by the same user twice, the other one is still open

OpenStudy (100101):

n1 + n2 = 2512 n2 = n1/384,000 n1 = 2512 - 2512/384,000